OpenStudy (anonymous):

Solve using the quadratic formula: 8x2 = 6x + 9

OpenStudy (anonymous):

i dont undestand how to use the quadratic formula with this. ?

OpenStudy (anonymous):

are you guys okay? it's taking a long time

OpenStudy (anonymous):

lol there probably showing all the work & explaining it

OpenStudy (anonymous):

8x2-6x-9=0 b=-6 a=8 c=-9 b2-4ac= 36+4*8*9= 324 sqrt(324)=18 so x= (6\[\pm\] 18)/(2*8)= 1.5,-0.75

OpenStudy (anonymous):

that is true. got to have some patience. OH we got a winner!!

OpenStudy (chaise):

8x^2 = 6x + 9 In order to solve quadratic equations using they MUST be equal to zero. We can rearrange the equation and effectively make it equal to zero by taking away 8x^2 from both sides. 0 = -8x^2+6x+9 This can be factorised to give: 0 = (-2x+3)(4x+3) x = -3/4 OR x = 3/2 I had trouble using the quadratic equation on this equation as the discriminate was under 0.

OpenStudy (anonymous):

and the hero finishes strong!